Tohban Report 2015-12-16
From RHESSI Wiki
|Start Date:||9 Dec 2015|
|End Date:||16 Dec 2015|
|List all reports|
Solar activity was low (at B-class level) through out the week although there were some C-class activities during the weekend. There are 9 ARs on disk. 4 ARs are expected to be rotated off the disk within 1 or 2 days.
How many GOES flares occurred?
Flares above B, C, M, X class were 6 1 0 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 2 1 0 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 2 0 0 0
There were RHESSI flares/GOES flares 183 / 7 over the time range 07-Dec-15 14-Dec-15
Decimation was active/vigorous for the full time period. Both front and rear detectors are off during SC night, except for 4 minutes on either side. The SSR is not emptying out.
The cold plates, 1 and 2 were at 132.5 and 131.1 K, on 2015-12-15.
No significantly large data gap.
GAPS IN APP_ID = 1 (VC1-SOH) WITH PACKET RATE LT 30
2015-12-11T19:35:00.000 -- 2015-12-11T19:50:00.000 900.00000
GAPS IN APP_ID = 154 (VC1-PMTRAS) WITH PACKET RATE LT 4
2015-12-11T19:40:00.000 -- 2015-12-11T19:50:00.000 600.00000
Early the week, we observed enhanced fast rates (and reduced livetime) in the rear segment of D1 and D7. Because rear-fast threshold is already 0xFF, we tried lowering HV. This helped D7, but there was no obvious change in the D1 monitor rates.
Dec 14, 18:19, lowered HV from 3600 to 3400 V.
Dec 09, 23:42, raised front fast threshold from 0x70 to 0xFF Dec 14, 18:19, lowered HV from 4100 to 3900 V
No improvement after power-cycling (The plan was to lower front slow threshold from 0xC0 to 0xB0 if the power-cycling could successfully reduce noise).
Dec 10, 16:42-18:19, power-cycling
HV was lowered in an attempt to reduce slow rates. The slow rates did decrease, but, when the events were turned on for a few minutes, the rate of write pointer motion increased significantly.
Dec 10, 21:38, lowered HV from 1500 V to 1000 V Dec 12, 00:33, events turned on for 3.5 min
No change made
D6 front slow threshold was lowered without significant increase of slow rates.
Dec 10, 19:58, lowered front slow threshold from 0x40 to 0x30 Dec 12, 00:30, lowered front slow threshold from 0x30 to 0x20
D7 front slow threshold was lowered to 0x10 but was brought back to 0x20 due to a significant increase of slow rates. D7 HV was lowered in attempt to reduce D7 rear fast rates (and to recover the livetime of the rear segment).
Dec 10, 21:38, lowered front slow threshold from 0x20 to 0x10 Dec 12, 00:33, lowered HV from 3500 to 3300 V Dec 12, 19:20, raised front slow threshold from 0x10 to 0x20
No change made
D9 is off
|HLAT Decimation||Rear decimation weight 6, no front decimation|
|Night time data (fronts)||+/- 4 minutes|
|Night time data (rears)||+/- 4 minutes|
|Require extra passes?||No|
|Requirement for moving pointer?||No|
|Attenuator operation||Working as expected|
|Detector problems?||See notes above.|